OpenStudy (anonymous):

Solve the following system of equations. Answer in an ordered pair. 2x+5y=16 -5x-2y=2 Help please!

OpenStudy (anonymous):

4x + 10y = 32 -25x - 10y = 10 -21x = 42 x = 42/-21

OpenStudy (anonymous):

and is y 0?

OpenStudy (anonymous):

Sorry you have to plug back in x correct? into one of the originals?

OpenStudy (anonymous):

x = -2 substitute 4(-2) +10y = 32 -8 +10y = 32 y =40/10 y=4

OpenStudy (anonymous):

yes ^^

OpenStudy (anonymous):

So the correct answer is (-2,4)?

OpenStudy (anonymous):

yupe

OpenStudy (anonymous):

Thank you